# The Software Diversity Card: A Framework for Reporting Diversity in Software Projects

### v1

@@ 7.2 Future work
In this section, we outline several directions for future research and development in reporting diversity-related aspects of software projects:
@@
Integrating diversity tracking to specific development process tasks: Right now we model the diversity of a project as a whole but we could follow a more fine-grained approach where diversity is also tracked at task level for deeper insights. For example, we could analyze whether a project has a diverse team but some tasks are only targeted by some team members, resulting in a low diverse approach for them (such as the project managers being mainly men). To generate these insights, it is essential to integrate Software Diversity Cards with software process definition tools and languages, such as SPEM [^24] SPEM specification homepage: https://www.omg.org/spec/SPEM/2.0/About-SPEM.
@@
Enhancing the diversity card with environmental information: Alongside diversity, there is a growing concern regarding the environmental aspects of software development. For instance, in the machine learning (ML) field, transparency about the environmental impact of AI training has become a significant topic. Initiatives like MyGreenLabel [^25] MyGreenLabel homepage: https://act.mygreenlab.org provide insights into the carbon footprint of trained ML models. Incorporating this environmental information into the Software Diversity Card framework could further contribute to the creation of ”Responsible Software.” Exploring methods to integrate and balance these two types of information would be a fruitful avenue for research.
@@ Acknowledgements
Joan Giner-Miguelez acknowledges his AI4S fellowship within the “Generación D” initiative by Red.es, Ministerio para la Transformación Digital y de la Función Pública, for talent attraction (C005/24-ED CV1), funded by NextGenerationEU through PRTR. This work is part of the projects TED2021-130331B-I00 and PID2023-147592OB-I00, and the research network RED2022-134647-T, all funded by MCIN/ AEI/10.13039/501100011033 and the Luxembourg National Research Fund (FNR) PEARL program, grant agreement 16544475.

### v2

@@ 8 Conclusions and future work
In the remainder of this section, we outline several directions for future research and development in reporting diversity-related aspects of software projects:
@@
Integrating diversity tracking to specific development process tasks: Right now, we model the diversity of a project as a whole, but we could follow a more fine-grained approach where diversity is also tracked at the task level for deeper insights. For example, we could analyze whether a project has a diverse team, but some tasks are only targeted by some team members, resulting in a less diverse approach for them (such as the project managers being mainly men). In addition, it would be interesting to account for the level of involvement of the individuals, as different levels of involvement could also raise bias and insights into diversity in a team. To generate these insights, it is essential to integrate Software Diversity Cards with software process definition tools and languages, such as SPEM [^22] SPEM specification homepage: https://www.omg.org/spec/SPEM/2.0/About-SPEM.
@@
Enhancing the diversity card with environmental information: Alongside diversity, there is a growing concern regarding the environmental aspects of software development. For instance, in the machine learning (ML) field, transparency about the environmental impact of AI training has become a significant topic. Initiatives like MyGreenLabel [^23] MyGreenLabel homepage: https://act.mygreenlab.org provide insights into the carbon footprint of trained ML models. Incorporating this environmental information into the Software Diversity Card framework could further contribute to the creation of “Responsible Software.” Exploring methods to integrate and balance these two types of information would be a fruitful avenue for research.
@@ Acknowledgements
Joan Giner-Miguelez acknowledges his AI4S fellowship within the “Generación D” initiative by Red.es, Ministerio para la Transformación Digital y de la Función Pública, for talent attraction (C005/24-ED CV1), funded by NextGenerationEU through PRTR. This work is part of the projects TED2021-130331B-I00 and PID2023-147592OB-I00, and the research network RED2022-134647-T, all funded by MCIN/ AEI/10.13039/501100011033; and the Luxembourg National Research Fund (FNR) PEARL program, grant agreement 16544475.